# arXiv:2602.23589v1[cs.CV] 27 Feb 2026

### v1

@@ 3 Method
Our approach comprises three steps: editable pseude diagram synthesis, rule-based hard contrastive data generation and structure-aware contrastive learning. The first step regenerates pseude diagram images and captions in editable format from original (uneditable format) diagram images (detailed in Sec. 3.1). In the subsequent step, a set of rule-based diagrammatic editing operations is applied to the editable pseude diagram data to generate hard positive and negative samples (detailed in Sec. 3.2). Lastly, model training is conducted via contrastive learning that attends the diagrammatic structual difference between the pseude diagrams and their hard samples, as outlined in Sec. 3.3. A visual summary of the process is provided in Fig. 1.
3.1 Editable pseude diagram synthesis
@@
The pseude diagram synthesis process is illustrated in Fig. 2. Let $R=\{r_{i}\}^{N^{R}}_{i=1}$ denote the set of original diagram images in raster-format. The process involves regenerating diagrams $D=\{d_{i,j}\}_{i=1:N^{R},j=1:N^{d}}=\{(v_{i,j},t_{i,j},c_{i,j})\}_{i=1:N^{R},j=1:N^{d}}$, where $v_{i,j}$ is the diagram image in an editable format such as vector graphics and $t_{i,j},c_{i,j}$ are the corresponding description and diagram definition codes (e.g., Mermaid code [12], referred to as “diagram codes” or simply “codes”), by the steps in Algorithm 12.
Algorithm 1 Editable pseude diagram synthesis
@@ 3.2 Rule-based hard contrastive data generation
A set of the image editing operations is applied to the pseude diagram data obtained by the steps in Sec. 3.1, as proposed in [17] but we modify the rules which outlined below:
@@ 4.2 Dataset Preparation
For the case of PSEUDE, only images in the flowchart dataset are used to constract the training data. The reconstructed (editable) data is generated by the process in Sec. 3.1. The node size $N^{n}$ is set to 3, as the same as the granulation process. We also use Tesseract OCR for the text extraction in the process. The total number of the reconstructed pseude diagrams is controlled by random sampling to match the size of the GRAN case. The flowchart images are rendered in the same manner as the GRAN case.
@@
PSEUDE (ours) - CLIP [16] 0.76667 0.93952 0.95645 0.84315 0.78253 0.93790 0.95591 0.85319
PSEUDE (ours) - SaCLIP [17] 0.77097 0.93952 0.95914 0.84542 0.78763 0.93871 0.95484 0.85532
@@
PSEUDE (ours) - CLIP [16] 0.35269 0.35269 0.35269 0.50511 0.76022 0.63644
PSEUDE (ours) - SaCLIP [17] 0.38817 0.38817 0.38817 0.70484 0.87204 0.79069
@@ 4.5 VQA Task Evaluation
PSEUDE - CLIP (ours) 0.558813 0.488994 0.536966
PSEUDE - SaCLIP (ours) 0.660028 0.612537 0.647963
@@ 4.6 Limitation
Whilst the experiments presented in this paper demonstrate the effectiveness of our method in improving performance on flowchart image-text matching and VQA tasks, a notable limitation regarding the dataset assumption remains. Our method heavily relies on the assumption of the similar appearance between diagrams in a target domain and the output of a diagram renderer. It is not fully applicable when the appearance of diagrams in a target domain is different from a style that the diagram renderer does not support. Furthermore, the quality of the pseude-diagrams is degraded when the type of the target diagram is not supported by the diagram renderer. This limites the range of application of our approach.

### v2

# arXiv:2602.23589v2[cs.CV] 02 Mar 2026

@@ 3 Method
Our approach comprises three steps: editable pseudo diagram synthesis, rule-based hard contrastive data generation and structure-aware contrastive learning. The first step regenerates pseudo diagram images and captions in editable format from original (uneditable format) diagram images (detailed in Sec. 3.1). In the subsequent step, a set of rule-based diagrammatic editing operations is applied to the editable pseudo diagram data to generate hard positive and negative samples (detailed in Sec. 3.2). Lastly, model training is conducted via contrastive learning that attends the diagrammatic structual difference between the pseudo diagrams and their hard samples, as outlined in Sec. 3.3. A visual summary of the process is provided in Fig. 1.
3.1 Editable pseudo diagram synthesis
@@
The pseudo diagram synthesis process is illustrated in Fig. 2. Let $R=\{r_{i}\}^{N^{R}}_{i=1}$ denote the set of original diagram images in raster-format. The process involves regenerating diagrams $D=\{d_{i,j}\}_{i=1:N^{R},j=1:N^{d}}=\{(v_{i,j},t_{i,j},c_{i,j})\}_{i=1:N^{R},j=1:N^{d}}$, where $v_{i,j}$ is the diagram image in an editable format such as vector graphics and $t_{i,j},c_{i,j}$ are the corresponding description and diagram definition codes (e.g., Mermaid code [12], referred to as “diagram codes” or simply “codes”), by the steps in LABEL:alg:pseuoe_synthesis2.
Algorithm 1 Editable pseudo diagram synthesis
@@ 3.2 Rule-based hard contrastive data generation
A set of the image editing operations is applied to the pseudo diagram data obtained by the steps in Sec. 3.1, as proposed in [17] but we modify the rules which outlined below:
@@ 4.2 Dataset Preparation
For the case of PSEUDO, only images in the flowchart dataset are used to constract the training data. The reconstructed (editable) data is generated by the process in Sec. 3.1. The node size $N^{n}$ is set to 3, as the same as the granulation process. We also use Tesseract OCR for the text extraction in the process. The total number of the reconstructed pseudo diagrams is controlled by random sampling to match the size of the GRAN case. The flowchart images are rendered in the same manner as the GRAN case.
@@
PSEUDO (ours) - CLIP [16] 0.76667 0.93952 0.95645 0.84315 0.78253 0.93790 0.95591 0.85319
PSEUDO (ours) - SaCLIP [17] 0.77097 0.93952 0.95914 0.84542 0.78763 0.93871 0.95484 0.85532
@@
PSEUDO (ours) - CLIP [16] 0.35269 0.35269 0.35269 0.50511 0.76022 0.63644
PSEUDO (ours) - SaCLIP [17] 0.38817 0.38817 0.38817 0.70484 0.87204 0.79069
@@ 4.5 VQA Task Evaluation
PSEUDO - CLIP (ours) 0.558813 0.488994 0.536966
PSEUDO - SaCLIP (ours) 0.660028 0.612537 0.647963
@@ 4.6 Limitation
Whilst the experiments presented in this paper demonstrate the effectiveness of our method in improving performance on flowchart image-text matching and VQA tasks, a notable limitation regarding the dataset assumption remains. Our method heavily relies on the assumption of the similar appearance between diagrams in a target domain and the output of a diagram renderer. It is not fully applicable when the appearance of diagrams in a target domain is different from a style that the diagram renderer does not support. Furthermore, the quality of the pseudo-diagrams is degraded when the type of the target diagram is not supported by the diagram renderer. This limites the range of application of our approach.